AMY CAN CLEAN THE ELEPHANT ENCLOSURE IN 4 HOURS, BUT IT TAKE JOEL 6 HOURS. HOW LONG WOULD IT TAKE THEM WORKING TOGETHER.....PLEASE HELP
Answer by nerdybill(7384)

AMY CAN CLEAN THE ELEPHANT ENCLOSURE IN 4 HOURS, BUT IT TAKE JOEL 6 HOURS. HOW LONG WOULD IT TAKE THEM WORKING TOGETHER.....PLEASE HELP
.
Let x = time (hours) it takes together
then
x(1/4 + 1/6) = 1
multiplying both sides by 12:
x(3 + 2) = 12
x(5) = 12
x = 12/5
x = 2.4 hours
or
x = 2 hours and 24 minutes